\begin{thebibliography}{43}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Antoun et~al.(2020)Antoun, Baly, and Hajj}]{arabert}
Wissam Antoun, Fady Baly, and Hazem Hajj. 2020.
\newblock \href {https://aclanthology.org/2020.osact-1.2} {{A}ra{BERT}:
  Transformer-based model for {A}rabic language understanding}.
\newblock In \emph{Proceedings of the 4th Workshop on Open-Source Arabic
  Corpora and Processing Tools, with a Shared Task on Offensive Language
  Detection}, pages 9--15, Marseille, France. European Language Resource
  Association.

\bibitem[{Artetxe et~al.(2020)Artetxe, Ruder, and Yogatama}]{Artetxe2019}
Mikel Artetxe, Sebastian Ruder, and Dani Yogatama. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.421} {On the
  cross-lingual transferability of monolingual representations}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 4623--4637, Online. Association for
  Computational Linguistics.

\bibitem[{Bhattacharjee et~al.(2021)Bhattacharjee, Hasan, Ahmad, Samin, Islam,
  Iqbal, Rahman, and Shahriyar}]{banglabert}
Abhik Bhattacharjee, Tahmid Hasan, Wasi~Uddin Ahmad, Kazi Samin, Md~Saiful
  Islam, Anindya Iqbal, M.~Sohel Rahman, and Rifat Shahriyar. 2021.
\newblock \href {https://doi.org/10.48550/ARXIV.2101.00204} {Banglabert:
  Language model pretraining and benchmarks for low-resource language
  understanding evaluation in bangla}.

\bibitem[{Blevins and Zettlemoyer(2022)}]{Blevins2022}
Terra Blevins and Luke Zettlemoyer. 2022.
\newblock \href {https://doi.org/10.48550/arxiv.2204.08110} {Language
  contamination explains the cross-lingual capabilities of english pretrained
  models}.

\bibitem[{Bochkarev et~al.(2015)Bochkarev, Shevlyakova, and
  Solovyev}]{bochkarev2015average}
Vladimir~V Bochkarev, Anna~V Shevlyakova, and Valery~D Solovyev. 2015.
\newblock The average word length dynamics as an indicator of cultural changes
  in society.
\newblock \emph{Social Evolution and History}, 14(2):153--175.

\bibitem[{Bommasani et~al.(2021)Bommasani, Hudson, Adeli, Altman, Arora, von
  Arx, Bernstein, Bohg, Bosselut, Brunskill, Brynjolfsson, Buch, Card,
  Castellon, Chatterji, Chen, Creel, Davis, Demszky, Donahue, Doumbouya,
  Durmus, Ermon, Etchemendy, Ethayarajh, Fei{-}Fei, Finn, Gale, Gillespie,
  Goel, Goodman, Grossman, Guha, Hashimoto, Henderson, Hewitt, Ho, Hong, Hsu,
  Huang, Icard, Jain, Jurafsky, Kalluri, Karamcheti, Keeling, Khani, Khattab,
  Koh, Krass, Krishna, Kuditipudi, and et~al.}]{foundation}
Rishi Bommasani, Drew~A. Hudson, Ehsan Adeli, Russ Altman, Simran Arora, Sydney
  von Arx, Michael~S. Bernstein, Jeannette Bohg, Antoine Bosselut, Emma
  Brunskill, Erik Brynjolfsson, Shyamal Buch, Dallas Card, Rodrigo Castellon,
  Niladri~S. Chatterji, Annie~S. Chen, Kathleen Creel, Jared~Quincy Davis,
  Dorottya Demszky, Chris Donahue, Moussa Doumbouya, Esin Durmus, Stefano
  Ermon, John Etchemendy, Kawin Ethayarajh, Li~Fei{-}Fei, Chelsea Finn, Trevor
  Gale, Lauren Gillespie, Karan Goel, Noah~D. Goodman, Shelby Grossman, Neel
  Guha, Tatsunori Hashimoto, Peter Henderson, John Hewitt, Daniel~E. Ho, Jenny
  Hong, Kyle Hsu, Jing Huang, Thomas Icard, Saahil Jain, Dan Jurafsky,
  Pratyusha Kalluri, Siddharth Karamcheti, Geoff Keeling, Fereshte Khani, Omar
  Khattab, Pang~Wei Koh, Mark~S. Krass, Ranjay Krishna, Rohith Kuditipudi, and
  et~al. 2021.
\newblock \href {http://arxiv.org/abs/2108.07258} {On the opportunities and
  risks of foundation models}.
\newblock \emph{CoRR}, abs/2108.07258.

\bibitem[{Carmo et~al.(2020)Carmo, Piau, Campiotti, Nogueira, and
  Lotufo}]{ptt5}
Diedre Carmo, Marcos Piau, Israel Campiotti, Rodrigo Nogueira, and Roberto
  Lotufo. 2020.
\newblock \href {https://doi.org/10.48550/ARXIV.2008.09144} {Ptt5: Pretraining
  and validating the t5 model on brazilian portuguese data}.

\bibitem[{Cañete et~al.(2020)Cañete, Chaperon, Fuentes, Ho, Kang, and
  Pérez}]{beto-model}
José Cañete, Gabriel Chaperon, Rodrigo Fuentes, Jou-Hui Ho, Hojin Kang, and
  Jorge Pérez. 2020.
\newblock Spanish pre-trained bert model and evaluation data.
\newblock In \emph{PML4DC at ICLR 2020}.

\bibitem[{Chan et~al.(2020)Chan, Schweter, and M{\"o}ller}]{germanbert}
Branden Chan, Stefan Schweter, and Timo M{\"o}ller. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.coling-main.598}
  {{G}erman{'}s next language model}.
\newblock In \emph{Proceedings of the 28th International Conference on
  Computational Linguistics}, pages 6788--6796, Barcelona, Spain (Online).
  International Committee on Computational Linguistics.

\bibitem[{Chiang and yi~Lee(2020)}]{pretrainwithouthuman}
Cheng-Han Chiang and Hung yi~Lee. 2020.
\newblock \href {https://doi.org/10.48550/arxiv.2012.11995} {Pre-training a
  language model without human language}.

\bibitem[{Clark et~al.(2020)Clark, Choi, Collins, Garrette, Kwiatkowski,
  Nikolaev, and Palomaki}]{tydiqa}
Jonathan~H. Clark, Eunsol Choi, Michael Collins, Dan Garrette, Tom Kwiatkowski,
  Vitaly Nikolaev, and Jennimaria Palomaki. 2020.
\newblock Tydi qa: A benchmark for information-seeking question answering in
  typologically diverse languages.
\newblock \emph{Transactions of the Association for Computational Linguistics}.

\bibitem[{Conneau et~al.(2019)Conneau, Khandelwal, Goyal, Chaudhary, Wenzek,
  Guzm{\'{a}}n, Grave, Ott, Zettlemoyer, and Stoyanov}]{Conneau2020}
Alexis Conneau, Kartikay Khandelwal, Naman Goyal, Vishrav Chaudhary, Guillaume
  Wenzek, Francisco Guzm{\'{a}}n, Edouard Grave, Myle Ott, Luke Zettlemoyer,
  and Veselin Stoyanov. 2019.
\newblock \href {http://arxiv.org/abs/1911.02116} {Unsupervised cross-lingual
  representation learning at scale}.
\newblock \emph{CoRR}, abs/1911.02116.

\bibitem[{Conneau et~al.(2018)Conneau, Rinott, Lample, Williams, Bowman,
  Schwenk, and Stoyanov}]{conneau2018xnli}
Alexis Conneau, Ruty Rinott, Guillaume Lample, Adina Williams, Samuel~R.
  Bowman, Holger Schwenk, and Veselin Stoyanov. 2018.
\newblock {"XNLI: Evaluating Cross-lingual Sentence Representations"}.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing}. Association for Computational Linguistics.

\bibitem[{Conneau et~al.(2020)Conneau, Wu, Li, Zettlemoyer, and
  Stoyanov}]{Wu2020}
Alexis Conneau, Shijie Wu, Haoran Li, Luke Zettlemoyer, and Veselin Stoyanov.
  2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.536} {Emerging
  cross-lingual structure in pretrained language models}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 6022--6034, Online. Association for
  Computational Linguistics.

\bibitem[{de~Souza et~al.(2021)de~Souza, Nogueira, and
  Lotufo}]{souza2021langagnostic}
Leandro~Rodrigues de~Souza, Rodrigo Nogueira, and Roberto Lotufo. 2021.
\newblock \href {https://doi.org/10.48550/ARXIV.2109.01942} {On the ability of
  monolingual models to learn language-agnostic representations}.

\bibitem[{Deshpande et~al.(2021)Deshpande, Talukdar, and
  Narasimhan}]{deshpande2021bert}
Ameet Deshpande, Partha Talukdar, and Karthik Narasimhan. 2021.
\newblock When is bert multilingual? isolating crucial ingredients for
  cross-lingual transfer.
\newblock \emph{arXiv preprint arXiv:2110.14782}.

\bibitem[{Devlin et~al.(2019)Devlin, Chang, Lee, and
  Toutanova}]{devlin-etal-2019-bert}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019.
\newblock \href {https://doi.org/10.18653/v1/N19-1423} {{BERT}: Pre-training of
  deep bidirectional transformers for language understanding}.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, pages 4171--4186,
  Minneapolis, Minnesota. Association for Computational Linguistics.

\bibitem[{Fujinuma et~al.(2022)Fujinuma, Boyd-Graber, and Kann}]{Fujinuma2022}
Yoshinari Fujinuma, Jordan Boyd-Graber, and Katharina Kann. 2022.
\newblock \href {https://doi.org/10.48550/arxiv.2203.10753} {Match the script,
  adapt if multilingual: Analyzing the effect of multilingual pretraining on
  cross-lingual transferability}.

\bibitem[{Heek et~al.(2020)Heek, Levskaya, Oliver, Ritter, Rondepierre,
  Steiner, and van {Z}ee}]{flax2020github}
Jonathan Heek, Anselm Levskaya, Avital Oliver, Marvin Ritter, Bertrand
  Rondepierre, Andreas Steiner, and Marc van {Z}ee. 2020.
\newblock \href {http://github.com/google/flax} {{F}lax: A neural network
  library and ecosystem for {JAX}}.

\bibitem[{Hu et~al.(2020)Hu, Ruder, Siddhant, Neubig, Firat, and
  Johnson}]{xtreme_paper}
Junjie Hu, Sebastian Ruder, Aditya Siddhant, Graham Neubig, Orhan Firat, and
  Melvin Johnson. 2020.
\newblock {XTREME}: A {M}assively {M}ultilingual {M}ulti-task {B}enchmark for
  {E}valuating {C}ross-lingual {G}eneralisation.
\newblock In \emph{International Conference on Machine Learning}, pages
  4411--4421. PMLR.

\bibitem[{Krishna et~al.(2021)Krishna, Bigham, and
  Lipton}]{krishna-etal-2021-pretraining-summarization}
Kundan Krishna, Jeffrey Bigham, and Zachary~C. Lipton. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.findings-emnlp.273} {Does
  pretraining for summarization require knowledge transfer?}
\newblock In \emph{Findings of the Association for Computational Linguistics:
  EMNLP 2021}, pages 3178--3189, Punta Cana, Dominican Republic. Association
  for Computational Linguistics.

\bibitem[{Kummervold(2022)}]{NorthT5}
Per~E Kummervold. 2022.
\newblock North {T5}.
\newblock \url{https://github.com/peregilk/north-t5}.

\bibitem[{Lee et~al.(2021)Lee, Yoon, Hwang, Joe, Min, and Gwon}]{korealbert}
Hyunjae Lee, Jaewoong Yoon, Bonggyu Hwang, Seongho Joe, Seungjai Min, and
  Youngjune Gwon. 2021.
\newblock \href {https://doi.org/10.48550/ARXIV.2101.11363} {Korealbert:
  Pretraining a lite bert model for korean language understanding}.

\bibitem[{Lu et~al.(2021)Lu, Abbeel, Grover, Mordatch, and Brain}]{Lu2021}
Kevin Lu, Pieter Abbeel, Aditya Grover, Igor Mordatch, and Google Brain. 2021.
\newblock \href {https://doi.org/10.48550/arxiv.2103.05247} {Pretrained
  transformers as universal computation engines}.

\bibitem[{Martin et~al.(2020)Martin, Muller, Ortiz~Su{\'a}rez, Dupont, Romary,
  de~la Clergerie, Seddah, and Sagot}]{camembert}
Louis Martin, Benjamin Muller, Pedro~Javier Ortiz~Su{\'a}rez, Yoann Dupont,
  Laurent Romary, {\'E}ric de~la Clergerie, Djam{\'e} Seddah, and Beno{\^\i}t
  Sagot. 2020.
\newblock \href {https://www.aclweb.org/anthology/2020.acl-main.645}
  {{C}amem{BERT}: a tasty {F}rench language model}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 7203--7219, Online. Association for
  Computational Linguistics.

\bibitem[{Nguyen and Tuan~Nguyen(2020)}]{phobert}
Dat~Quoc Nguyen and Anh Tuan~Nguyen. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.findings-emnlp.92}
  {{P}ho{BERT}: Pre-trained language models for {V}ietnamese}.
\newblock In \emph{Findings of the Association for Computational Linguistics:
  EMNLP 2020}, pages 1037--1042, Online. Association for Computational
  Linguistics.

\bibitem[{Oladipo et~al.(2022)Oladipo, Ogundepo, Ogueji, and Lin}]{jimmylin}
Akintunde Oladipo, Odunayo Ogundepo, Kelechi Ogueji, and Jimmy Lin. 2022.
\newblock \href {https://openreview.net/forum?id=HOZmF9MV8Wc} {An exploration
  of vocabulary size and transfer effects in multilingual language models for
  african languages}.
\newblock In \emph{3rd Workshop on African Natural Language Processing}.

\bibitem[{Papadimitriou and
  Jurafsky(2020)}]{papadimitriou-jurafsky-2020-learning}
Isabel Papadimitriou and Dan Jurafsky. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.emnlp-main.554} {{L}earning
  {M}usic {H}elps {Y}ou {R}ead: {U}sing transfer to study linguistic structure
  in language models}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 6829--6839, Online. Association
  for Computational Linguistics.

\bibitem[{Pires et~al.(2019)Pires, Schlinger, and Garrette}]{Pires2019}
Telmo Pires, Eva Schlinger, and Dan Garrette. 2019.
\newblock \href {http://arxiv.org/abs/1906.01502} {{How multilingual is
  Multilingual BERT?}}
\newblock \emph{ACL 2019 - 57th Annual Meeting of the Association for
  Computational Linguistics, Proceedings of the Conference}, pages 4996--5001.

\bibitem[{Radford et~al.(2019)Radford, Wu, Child, Luan, Amodei, Sutskever
  et~al.}]{radford2019language}
Alec Radford, Jeffrey Wu, Rewon Child, David Luan, Dario Amodei, Ilya
  Sutskever, et~al. 2019.
\newblock Language models are unsupervised multitask learners.
\newblock \emph{OpenAI blog}, 1(8):9.

\bibitem[{Raffel et~al.(2020)Raffel, Shazeer, Roberts, Lee, Narang, Matena,
  Zhou, Li, and Liu}]{t5}
Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael
  Matena, Yanqi Zhou, Wei Li, and Peter~J. Liu. 2020.
\newblock \href {http://jmlr.org/papers/v21/20-074.html} {Exploring the limits
  of transfer learning with a unified text-to-text transformer}.
\newblock \emph{Journal of Machine Learning Research}, 21(140):1--67.

\bibitem[{Rajpurkar et~al.(2016)Rajpurkar, Zhang, Lopyrev, and Liang}]{squad}
Pranav Rajpurkar, Jian Zhang, Konstantin Lopyrev, and Percy Liang. 2016.
\newblock \href {https://doi.org/10.48550/ARXIV.1606.05250} {{SQuAD: 100,000+
  Questions for Machine Comprehension of Text}}.

\bibitem[{Real et~al.(2020)Real, Fonseca, and Gon{\c{c}}alo~Oliveira}]{assin2}
Livy Real, Erick Fonseca, and Hugo Gon{\c{c}}alo~Oliveira. 2020.
\newblock The assin 2 shared task: A quick overview.
\newblock In \emph{Computational Processing of the Portuguese Language}, pages
  406--412, Cham. Springer International Publishing.

\bibitem[{Ri and Tsuruoka(2022)}]{pretrainingartificial}
Ryokan Ri and Yoshimasa Tsuruoka. 2022.
\newblock \href {https://doi.org/10.48550/ARXIV.2203.10326} {Pretraining with
  artificial language: Studying transferable knowledge in language models}.

\bibitem[{Rust et~al.(2021)Rust, Pfeiffer, Vuli{\'c}, Ruder, and
  Gurevych}]{rust-etal-2021-good}
Phillip Rust, Jonas Pfeiffer, Ivan Vuli{\'c}, Sebastian Ruder, and Iryna
  Gurevych. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.acl-long.243} {How good is
  your tokenizer? on the monolingual performance of multilingual language
  models}.
\newblock In \emph{Proceedings of the 59th Annual Meeting of the Association
  for Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)}, pages 3118--3135,
  Online. Association for Computational Linguistics.

\bibitem[{Sarti and Nissim(2022)}]{it5}
Gabriele Sarti and Malvina Nissim. 2022.
\newblock \href {https://doi.org/10.48550/ARXIV.2203.03759} {It5: Large-scale
  text-to-text pretraining for italian language understanding and generation}.

\bibitem[{{Sayama} et~al.(2019){Sayama}, {Araujo}, and
  {Fernandes}}]{faquad2019}
H.~F. {Sayama}, A.~V. {Araujo}, and E.~R. {Fernandes}. 2019.
\newblock {FaQuAD: Reading Comprehension Dataset in the Domain of Brazilian
  Higher Education}.
\newblock In \emph{2019 8th Brazilian Conference on Intelligent Systems
  (BRACIS)}, pages 443--448.

\bibitem[{Shazeer and Stern(2018)}]{adafactor}
Noam Shazeer and Mitchell Stern. 2018.
\newblock \href {http://arxiv.org/abs/1804.04235} {Adafactor: Adaptive learning
  rates with sublinear memory cost}.

\bibitem[{Souza et~al.(2020)Souza, Nogueira, and Lotufo}]{bertimbau}
Fábio Souza, Rodrigo Nogueira, and Roberto Lotufo. 2020.
\newblock \href {https://doi.org/10.1007/978-3-030-61377-8_28/FIGURES/2}
  {Bertimbau: Pretrained bert models for brazilian portuguese}.
\newblock \emph{Lecture Notes in Computer Science (including subseries Lecture
  Notes in Artificial Intelligence and Lecture Notes in Bioinformatics)}, 12319
  LNAI:403--417.

\bibitem[{Vania et~al.(2021)Vania, Htut, Huang, Mungra, Pang, Phang, Liu, Cho,
  and Bowman}]{Vania2021}
Clara Vania, Phu~Mon Htut, William Huang, Dhara Mungra, Richard~Yuanzhe Pang,
  Jason Phang, Haokun Liu, Kyunghyun Cho, and Samuel~R. Bowman. 2021.
\newblock \href {https://doi.org/10.18653/V1/2021.ACL-LONG.92} {Comparing test
  sets with item response theory}.
\newblock \emph{ACL-IJCNLP 2021 - 59th Annual Meeting of the Association for
  Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing, Proceedings of the Conference}, pages
  1141--1158.

\bibitem[{Xue et~al.(2021{\natexlab{a}})Xue, Barua, Constant, Al-Rfou, Narang,
  Kale, Roberts, and Raffel}]{xue2021byt5}
Linting Xue, Aditya Barua, Noah Constant, Rami Al-Rfou, Sharan Narang, Mihir
  Kale, Adam Roberts, and Colin Raffel. 2021{\natexlab{a}}.
\newblock \href {http://arxiv.org/abs/2105.13626} {Byt5: Towards a token-free
  future with pre-trained byte-to-byte models}.

\bibitem[{Xue et~al.(2021{\natexlab{b}})Xue, Constant, Roberts, Kale, Al-Rfou,
  Siddhant, Barua, and Raffel}]{xue-etal-2021-mt5}
Linting Xue, Noah Constant, Adam Roberts, Mihir Kale, Rami Al-Rfou, Aditya
  Siddhant, Aditya Barua, and Colin Raffel. 2021{\natexlab{b}}.
\newblock \href {https://doi.org/10.18653/v1/2021.naacl-main.41} {m{T}5: A
  massively multilingual pre-trained text-to-text transformer}.
\newblock In \emph{Proceedings of the 2021 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies}, pages 483--498, Online. Association for Computational
  Linguistics.

\bibitem[{Zhang et~al.(2020)Zhang, Warstadt, Li, and Bowman}]{bowman}
Yian Zhang, Alex Warstadt, Haau-Sing Li, and Samuel~R. Bowman. 2020.
\newblock \href {https://doi.org/10.48550/ARXIV.2011.04946} {When do you need
  billions of words of pretraining data?}

\end{thebibliography}

\pdfoutput=1
\documentclass[11pt]{article}

\usepackage{acl}

\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}
\usepackage[utf8]{inputenc}

\usepackage{microtype}
\usepackage{times}
\usepackage{url}
\usepackage{latexsym}
\usepackage{booktabs}

\usepackage{graphicx}

\usepackage{todonotes}
\usepackage{hyperref}

\title{MonoByte: A Pool of Monolingual Byte-level Language Models}

\author{Hugo Abonizio \\
  FEEC, UNICAMP, Brazil \\
  NeuralMind, Brazil \\
  {\tt hugo.abonizio@gmail.com} \\
  \And
  Leandro Rodrigues de Souza \\
  FEEC, UNICAMP, Brazil \\
  {\tt l231250@g.unicamp.br} \\
  \AND
  Roberto Lotufo \\
  FEEC, UNICAMP, Brazil \\
  NeuralMind, Brazil \\
  {\tt lotufo@unicamp.br } \\
  \And
  Rodrigo Nogueira \\
  FEEC, UNICAMP, Brazil \\
  NeuralMind, Brazil \\
  {\tt rfn@unicamp.br} \\
}

\date{}

\begin{document}
\maketitle
\begin{abstract}
  The zero-shot cross-lingual ability of models pretrained on multilingual and even monolingual corpora has spurred many hypotheses to explain this intriguing empirical result. However, due to the costs of pretraining, most research uses public models whose pretraining methodology, such as the choice of tokenization, corpus size, and computational budget, might differ drastically. When researchers pretrain their own models, they often do so under a constrained budget, and the resulting models might underperform significantly compared to SOTA models.
  These experimental differences led to various inconsistent conclusions about the nature of the cross-lingual ability of these models.
  To help further research on the topic, we released 10 monolingual byte-level models\footnote{\url{https://huggingface.co/monobyte}} rigorously pretrained under the same configuration with a large compute budget (equivalent to 420 days on a V100) and corpora that are 4 times larger than the original BERT's. Because they are tokenizer-free, the problem of unseen token embeddings is eliminated, thus allowing researchers to try a wider range of cross-lingual experiments in languages with different scripts.
  Additionally, we release two models pretrained on non-natural language texts that can be used in sanity-check experiments.
  Experiments on QA and NLI tasks show that our monolingual models achieve competitive performance to the multilingual one, and hence can be served to strengthen our understanding of cross-lingual transferability in language models.
\end{abstract}

\section{Introduction}

Shortly after the publication of BERT~\cite{devlin-etal-2019-bert}, researchers showed that multilingual models pretrained on the masked language modeling objective can achieve remarkable zero-shot cross-lingual performance on various NLP tasks (i.e., a multilingual model finetuned on a high-resource language and directly evaluated in other languages)~\cite{Conneau2020, xtreme_paper}. 

% For example, a model pretrained on a multilingual corpus and finetuned on a task that contains only English texts can perform well on that same task but on texts in other languages~\cite{Pires2019, xtreme_paper}.

These empirical results triggered a wave of research that aimed to explain this behavior. \citet{Pires2019} raised the ``anchor tokens'' hypothesis, i.e., that tokens shared between two languages act as a reference point so models can learn similar concepts. \citet{Artetxe2019} and \citet{Wu2020} questioned these findings and pointed out that, even without a shared vocabulary, models achieve great cross-lingual transfer performance by leveraging shared parameters only.

Perhaps even more surprising is the performance of \textit{monolingual} models in a cross-lingual setting. For instance, models pretrained and finetuned only on English can perform well in French tasks~\cite{jimmylin}. \citet{souza2021langagnostic} also show that monolingual models finetuned on a foreign language (i.e., a language that differs from pretraining) achieve comparable results to models finetuned on their language.

Recent work investigates properties of the pretraining corpora that contribute to the model's performance on natural language tasks. Models are pretrained on non-natural language corpora, such as code, music, proteins, and artificial languages and their performances are compared with natural language models~\cite{papadimitriou-jurafsky-2020-learning, pretrainwithouthuman, Lu2021, pretrainingartificial}. Evidence shows that pretraining on a corpus that contains artificial recursive or hierarchical structure between tokens results in a similar performance when compared to models pretrained on a natural language.

However, an outstanding problem is that these experiments are made either 1) using models pretrained by different research groups, whose pretraining configuration differs widely, or 2) using models pretrained under a compute budget or pretraining corpus that are orders of magnitude smaller than the ones used by SOTA models. For instance, \citet{pretrainwithouthuman} pretrain their models on a dataset 200 times smaller than BERT's. This is problematic because certain skills are learned only when corpus size and training budget are large enough~\cite{bowman}.

These differences in the pretraining methodology and undertrained models make it difficult to draw a conclusion about the nature of the cross-lingual ability of such models. Additionally, since most models use subword tokenization, it is difficult to experiment with languages that do not have a significant subword vocabulary overlap. For example, if a model is pretrained and finetuned on English, it cannot be tested in Chinese because there is very little token overlap in their vocabularies, and hence, Chinese token embeddings are not learned.
This is a problem even for languages with the same script. For example, many Spanish embeddings are not learned in an English-only pretraining, and hence it is difficult to tell whether a model's inability to learn cross-lingual representations is due only to tokenization issues~\cite{rust-etal-2021-good}.

Following \citet{foundation}, who advocate for the release of pretraining models, corpus, and script as a way to strengthen the field, we release byte-level models pretrained on large corpora from the same domain and using exactly the same training setup and compute budget.
Because they rely only on bytes to represent strings, they can be used to compare languages that use different scripts.

Each model takes approximately 210 hours of pretraining, resulting in more than three months of TPU computing. The models are available at \href{https://huggingface.co/monobyte}{https://huggingface.co/monobyte} and the code used for finetuning can be found at \href{https://github.com/lersouza/lang-agnostic}{https://github.com/lersouza/lang-agnostic}.

\section{Related work}

% \medskip
% \noindent\textbf{Monolingual models}. 
In the last years, many monolingual versions of BERT~\cite{devlin-etal-2019-bert, bertimbau, germanbert, camembert, arabert, korealbert, phobert, beto-model, banglabert} and T5~\cite{t5, ptt5, it5} have been released. The authors often claim that these versions outperform multilingual models in their languages. \citet{rust-etal-2021-good} showed that the tokenizer plays a critical role in achieving those results. More recently, \citet{xue2021byt5} released the ByT5 model, which would overcome this issue by leveraging a byte-level vocabulary. This model is pretrained on several languages and achieves great results when compared to mT5 \cite{xue-etal-2021-mt5}. However, no monolingual version of ByT5 has been released, which makes it difficult to conduct further investigation on the model's pretraining and cross-lingual performance that is not impacted by the tokenization in different languages.

% \medskip
% \noindent\textbf{Model pretraining research}.
Current research on the properties of pretraining often relies on monolingual models or models pretrained on artificial languages. Researchers often rely on models released by other groups~\cite{souza2021langagnostic} or pretrain models in a very controlled setup~\cite{Artetxe2019, papadimitriou-jurafsky-2020-learning, Fujinuma2022}. In both cases, it is not clear if the results suffer from these different methodologies or undertrained models. \citet{Blevins2022}, for instance, suggests that language contamination (i.e., a monolingual corpus composed of sentences in more languages) is the reason for monolingual zero-shot effectiveness. The same happens with models pretrained on non-natural language corpora~\cite{pretrainwithouthuman, Lu2021, pretrainingartificial}, where the corpus size differs from the ones trained on natural language, which might affect the conclusions drawn from the experiments. We hope to bridge this gap by releasing the set of monolingual models presented in this paper.

\section{Pretraining}

To train our monolingual models, we used mC4~\cite{xue-etal-2021-mt5}, the same corpus used to train mT5 and ByT5, which comprises 101 natural languages extracted from the Common Crawl web scrape. For each model, we selected only the documents written in the specific language, which was originally identified using \texttt{cld3} by \citet{xue-etal-2021-mt5}. The mC4 is a large corpus with different language distributions, with some being more represented than others. Thus, we trimmed each pretraining corpus in approximately 65 billion UTF-8 bytes for all languages.
The only exception was Bengali, which was trimmed on 32 billion bytes, corresponding to its total size.
% The only two exceptions were Bengali and Swahili, which were trimmed on 32 billion and 3 billion bytes, respectively, which are their total sizes.

The pretraining was conducted on a TPU VM v3-8 using Flax~\cite{flax2020github} library and the pretraining script for T5-like span-masked language modeling available on HuggingFace Transformers library.\footnote{\url{https://github.com/huggingface/transformers/blob/main/examples/flax/language-modeling/run_t5_mlm_flax.py}} We chose the smaller architecture of ByT5, with 300 million parameters, and used similar hyper-parameters as reported on ByT5~\cite{xue2021byt5} and mT5~\cite{xue-etal-2021-mt5} papers. We did not experiment with larger models due to their computational cost. We set the sequence length to 1024 tokens (UTF-8 bytes) and train for 1 million steps with a batch of $2^{16}$ bytes, resulting in approximately 65 billion bytes. Since we could only fit a batch of 8 examples per device, we used gradient accumulation of 128 steps to achieve a larger batch size. The learning rate starts at $10^{-3}$ and decays linearly throughout the pretraining.

In comparison with the original ByT5, our models are pretrained on a 16 times smaller corpus. However, our pretraining corpus is 4 times larger than BERT's, which was trained on 3,300M words~\cite{devlin-etal-2019-bert} --- i.e., approximately 16 billion characters, considering an average of 5 characters per word in English~\cite{bochkarev2015average}. In comparison with GPT-2, our corpus is 1.6 times larger~\cite{radford2019language}.

% Comparing with other monolingual models research, \citet{rust-etal-2021-good} pretrained for 900,000 steps using a batch size of 64 and a sequence length of 128 and the remaining 100,000 steps with 512, resulting in approximately 10 billion tokens. Thus, their experiments using the mBERT tokenizer, which has an average byte length per token of 7.3, resulted in 77GB of pretraining. \citet{pretrainingartificial} generated 12.8M sentences based on the average English sentence length, which is approximately 20 words~\cite{flesch1979write}, resulting in 1.2 billion characters or 1.2GB of artificial text.
Compared with other monolingual models research, \citet{rust-etal-2021-good} pretrained on a corpus of approximately 77GB, according to the reported number of steps and tokens per batch. \citet{pretrainingartificial} pretrained on a corpus of 1.2GB of artificially generated text.
A more extensive comparison to other work is presented in Table~\ref{tab:corpus-sizes} and demonstrates that our pretraining is comparable to SOTA monolingual models and much larger than corpora used in the majority of monolingual studies.

% However, each model takes approximately 210 hours of pretraining, resulting in more than 2,100 hours of TPU, or USD 18,480 (USD 8.8/hour). Although our monolingual models may not achieve the best performance when compared to their multilingual counterparts, our objective is to provide monolingual tokenizer-free language models to the research community...

% CamemBERT 138GB
% FinBERT 4.9 billion tokens
% BERTimbau 17GB
% X IndoBERT 220M words
% FlauBERT 71GB
% AraBERT 24GB
% BETO 3 bilion words (https://users.dcc.uchile.cl/~jperez/papers/pml4dc2020.pdf)
% KoreALBERT 43GB (4.4B words 18B characters)
% AraELECTRA 77GB or 8.8B words
% BanglaBERT 35GB (bangla ou bengali)
% PhoBERT Vietnamese 20GB

% \subsection{Languages}

\begin{table}
    %\small
    \centering\centering\resizebox{0.50\textwidth}{!}{
    \begin{tabular}{lp{1.8cm}r}
    \toprule
    \textbf{Model name} & \textbf{Language} & \textbf{Size} \\
    \midrule
    PTT5~\cite{ptt5}                  & pt & 15 \\
    BERT~\cite{devlin-etal-2019-bert} & en & 16 \\
    BERTimbau~\cite{bertimbau}        & pt & 17 \\
    PhoBERT~\cite{phobert}            & vi & 20 \\
    AraBERT~\cite{arabert}               & ar & 24 \\
    BanglaBERT~\cite{banglabert}        & bn & 35 \\
    KoreALBERT~\cite{korealbert}            & ko & 43 \\
    North ByT5~\cite{NorthT5}                   & no & 45 \\
    %BETO                  & es & \\
    CamemBERT~\cite{camembert}             & fr & 138 \\
    IT5~\cite{it5}                   & it & 215 \\
    \midrule
    \citealp{pretrainwithouthuman}  & en, artificial                    & 0.08 \\
    \citealp{papadimitriou-jurafsky-2020-learning}             & many, artif.                       & 0.5 \\
    \citealp{deshpande2021bert}     & many         & 0.5 \\
    \citealp{pretrainingartificial} & many, artif. & 1.2 \\
    \citealp{rust-etal-2021-good}              & many   & 77 \\
    \midrule
    MonoByte (ours, each model) & many & 65 \\
    \bottomrule
    \end{tabular}
    }
    \caption{Comparison of pretraining corpus sizes in GB of monolingual models. Models which does not explicitly report the corpus size were estimated based on the reported amount of tokens.}
    \label{tab:corpus-sizes}
\end{table}

\begin{table*}[h]
    \centering
    \small
    \begin{tabular}{lcccccccc}
        \toprule
        &                       \textbf{de} & \textbf{en} & \textbf{es} & \textbf{pt} & \textbf{ru} & \textbf{vi} & \textbf{zh} & \textbf{avg.} \\
        \midrule
        \textbf{Nonsense} & 33.33 & 33.55 & 33.27 & 50.00 & 33.33 & 33.57 & 34.19 & 35.89 \\
        \textbf{Hierarchical} & 66.01 & 69.32 & 68.04 & 65.65 & 62.12 & 64.49 & 62.83 & 65.49 \\
        \midrule
        \textbf{Monolingual}    &  70.85  & 76.76 &  75.07  &  83.40    & 67.39     & 70.37     &  69.91 & 73.39  \\
        \textbf{Multilingual}    & 76.02 &  80.55 &  78.39  &  67.73    &   73.58   &  73.01     &  73.39 & 74.67  \\
        \midrule
        $\Delta (\small{\textit{multi - mono}})$ & 5.16 & 3.80 & 3.32 & -15.67 & 6.19 &	2.64 &	3.48 & 1.28 \\
        \bottomrule
    \end{tabular}
    \caption{Results (accuracy) for Natural Language Inference in each language.}
    \label{table:results-nli}
    
\end{table*}

\begin{table*}
    \centering
    \small
    \begin{tabular}{lccccccc}
        \toprule
        &                       \textbf{ar} & \textbf{bn} & \textbf{en} & \textbf{ko} & \textbf{pt} & \textbf{ru} & \textbf{avg.} \\
        \midrule
                \textbf{Nonsense} & 0.22 & 0.00 & 0.45 & 0.00 & 0.26 & 0.27 & 0.20 \\
        \textbf{Hierarchical} & 67.23 & 0.38 & 28.03 & 9.99 & 3.63 & 32.85 & 23.68 \\
        \midrule
        \textbf{Monolingual}    & 78.17    & 53.89     &  60.15  & 51.17  & 64.88     & 58.26 & 61.09   \\
        \textbf{Multilingual}   & 81.81    & 70.28     &  68.51  & 54.70  & 42.37   & 72.97  & 65.11 \\ \midrule
        $\Delta (\small{\textit{multi - mono}})$ & 3.65 & 16.39 &	8.36 & 3.54 & -22.51 &	14.71 & 4.02 \\
        \bottomrule
    \end{tabular}
    \caption{Results (F1 Score) for Question Answering in each language.}
    \label{table:results-qa}
    
\end{table*}

% \subsection{Baselines}
% \noindent\textbf{Baselines}.
To sanity-check and validate the performance of our pretrained models, we compared their results with different baselines proposed in the literature. The first synthetic pretraining strategy, called nesting parentheses, was proposed by~\citet{papadimitriou-jurafsky-2020-learning}, which recursively generates random symbols respecting a hierarchical structure. This was proposed to evaluate how this recursion structure of the pretraining corpus transfers into modeling real languages. Their work showed that the models pretrained on the synthetic hierarchical corpus were able to predict human language far better than other baselines (e.g., random, music, and code), with a perplexity comparable to real languages.

% The first and more straightforward sanity check it to compare with models without pretraining, i.e., random initialized model weights. However, previous work have shown that models without pretraining tend to perform poorly on downstream tasks~\cite{CITAR}. Therefore, we also evaluated more strong pretraining strategies that use synthetic examples.

We used the code released by the authors, with the vocabulary being the 50,000 most frequent words based on the first one million documents of the Spanish mC4 corpus, and their sampling probability based on the Zipf distribution over the same Spanish corpus.\footnote{\url{https://github.com/toizzy/tilt-transfer/}} We generated examples until reaching the same size as our natural languages corpora sizes to make a fair comparison.

The second pretraining baseline was proposed by~\citet{krishna-etal-2021-pretraining-summarization}, dubbed nonsense, where the authors explored the knowledge transfer hypothesis of the downstream performance improvement of transfer learning. The authors pretrained on documents consisting of random n-grams and achieve similar performance on summarization over different datasets when compared to pretraining on real language. We generated examples using the original code\footnote{\url{https://github.com/acmi-lab/pretraining-with-nonsense/} \nopagebreak} to match the same corpus size of our natural language models.

\section{Experiments}

Our models are evaluated on two downstream tasks: Natural Language Inference (NLI) and Question Answering (QA). We use a similar setting as the \textit{In-language model}, described by \citet{xtreme_paper}, where the model is finetuned and evaluated in the subset of the task that corresponds to its pretraining language. The following subsections provide details about the datasets used and the results. The monolingual models are compared to the ByT5 Small checkpoint. The latter is expected to perform better since it is pretrained on a much larger set of a more diverse corpus, as investigated by \citet{Fujinuma2022}. For additional information  on the finetuning procedure, please refer to Appendix \ref{appendix:finetuning_details}.

% \subsection{Natural Language Inference}
\medskip
\noindent\textbf{Natural Language Inference}.
The XNLI\footnote{We use the machine-translated version of XNLI for training in different languages. Available at \href{https://dl.fbaipublicfiles.com/XNLI/XNLI-MT-1.0.zip}{https://dl.fbaipublicfiles.com/XNLI/XNLI-MT-1.0.zip}} dataset~\cite{conneau2018xnli} is used for finetuning and evaluating our German, English, Spanish, Russian, Vietnamese, and Chinese models. For the Portuguese model, we select the Recognizing Textual Entailment (RTE) task of the ASSIN2 dataset~\cite{assin2}. Performance is measured by using the accuracy metric.

% \medskip
Results are reported in Table \ref{table:results-nli}. Compared to the multilingual ByT5, our models achieve competitive performance. The difference in accuracy is about 1.28 on average. The results in Russian represent the largest gap (6.19 points), while those in Vietnamese represent the smallest (2.64 points). Nonsense achieves a near random performance, while Hierarchical approaches the models pretrained on natural language, as also evidenced by \citet{papadimitriou-jurafsky-2020-learning} and \citet{pretrainwithouthuman}.

% \subsection{Question Answering}
\medskip
\noindent\textbf{Question Answering}.
The gold passage version of the TydiQA~\cite{tydiqa} dataset (\textbf{TydiQA-GoldP}) is selected for finetuning and evaluating the Arabic, Bengali, English, Korean and Russian models. For finetuning the Portuguese model, we use the FaQuAD~\cite{faquad2019}. Performance is measured by the F1 Score metric as described by \citet{squad}.

Results are reported in Table~\ref{table:results-qa}. For Portuguese, our model outperforms the Multilingual checkpoint by 22.51 points, while, for Bengali, it stays 16.39 points behind. Our monolingual models stay, on average, 4 points behind the multilingual model released by Google. Looking at the difference between monolingual and multilingual models for every language, we see a higher variation. Nonsense and Hierarchical results are very distant. Question answering is a more difficult task~\cite{Vania2021} and a good benchmark for model quality. We hypothesize that this difference in task complexity requires more from the model. Pretraining with token structure emphasis only may not provide the required knowledge for good performance.

\section{Conclusion}

In this work, we introduced a pool of 10 tokenizer-free language models pretrained on large monolingual corpora. We demonstrated that our monolingual models can achieve competitive results to the multilingual ByT5, although having a smaller and less diverse pretraining --- single language, compared to all languages concatenated.

The main goal of our work was to release a set of monolingual models that carefully follow the same pretraining methodology to support the research of monolingual language models, pretraining properties, and cross-linguality. We hope that by releasing our models, we bridge this gap of having a more controlled and reproducible setup for rigorous experiments.

\section*{Acknowledgments}
This research was partially funded by grants 2020/09753-5 and 2022/01640-2 from Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP).
We also would like to thank Google Cloud for credits to support this work.

% include your own bib file like this:
\bibliographystyle{acl_natbib}
\bibliography{main}

\clearpage
\appendix

\section{Finetuning Details}
\label{appendix:finetuning_details}

The experiments were carried out in a single NVIDIA A100 80GB GPU. We use the Adafactor optimizer~\cite{adafactor} with a constant learning rate of $10^{-4}$. The models were finetuned using three different seeds, and we report an average of the results. We have chosen hyperparameters based on preliminary experiments with the ByT5 Small checkpoint\footnote{Available at: \url{https://huggingface.co/google/byt5-small}} on both XNLI and TydiQA datasets. We do not perform any exhaustive hyperparameter search, and use the same settings (per-task) for all models.

\medskip
\noindent\textbf{Natural Language Inference}. For all NLI experiments, we use a batch size of 16, accumulating gradients for 4 steps. The maximum input length was 1024, trimmed by batch. We train the model to output the class identifier (a number). We train for 10 epochs and evaluate every 0.2 epoch. We also perform early stopping with patience of 5 evaluations and select the best model on the validation set of each task. The results are reported against the test set of each dataset.

\medskip
\noindent\textbf{Question Answering}. For QA, the selected batch size is 6, accumulating gradients for 4 steps. The maximum input length was 2048 (question and context concatenated), trimmed by batch. We train the model to output an answer to the question with a maximum length of 768 bytes. We train for 10 epochs and evaluate at the end of each epoch. We also perform early stopping with patience of 3 and select the best model on the validation set of each task. 

\end{document}